Astronomy Report – December 28th

Name: Cesare Guariniello

Crew: 218

Date: 28-12-2019

MUSK OBSERVATORY

Practiced using the telescope both with the eyepiece and with the new camera. Let some of the crewmembers observe the sun. Practiced using the laptop software with the new camera.

Solar Features Observed: no visible sun spots, no visible prominences

Images submitted with this report: none

Problems encountered: still a little hard to have good visible solar features on the live stream from the CCD camera. Will work on imaging in the next few days.

Sol Summary – December 28th

Sol: 06

Summary Title: Still… Still… Waitin’ On A Sunny Day

Author’s name: Pat Pesa

Mission Status: Crew is excited for the good weather to come!

Sol Activity Summary: Today some progress was made towards making the crew video reports and training guides. Also some of the crew got to spend some time in the observatory to observe Sun granulation. Work also continued with the Engineer’s Indoor Climate-Comfort project.

Look ahead plan: Hoping for an afternoon EVA for tomorrow after more melting occurs.

Anomalies in work: None

Weather: Mostly Sunny, around 30 degrees

Crew physical status: Healthy

EVA: none

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, science report

Support Requested: none

Pat Pesa
Geologist, MDRS 218

Commander Report – December 28th

Crew 218 Commander Report 28-12-2019

Sol 6 – What day is it?

This morning we finally woke up to some sunlight. Despite cold temperatures and a few hours of fight with the clouds, the sun began melting some of the snow. At times it is hard to figure out what day it is, since we are isolated from most that happens on Earth. In the meanwhile, we continued our indoor activities with some video projects after stretching and breakfast. In the afternoon, we swept the icy snow from the tunnels between the buildings, and we managed to do some observations at the solar observatory. No features, but with the help of the crew astronomer some of the team members could observe the sun both with an eyepiece and with the CCD camera. We are hoping that snow will melt enough to give us time for one or two short EVAs at a later time tomorrow. Sun and cookies sure help with the mission!

Cesare Guariniello, Commander

style=”font-size:12pt”> Report – style=”font-size:12pt”> 27th

Journalist Report

Ben Durkee, Crew 218 Journalist

Sol 05

Today we awoke to find our pump to transfer water from the static tank to the loft tank was non-operational! The static tank is our outdoor water reserve containing all of our liquid life for the entire mission, whereas the loft tank is the tank in the Hab that supplies all of our utilities with water directly. In short, this anomaly meant we had to strictly ration our water usage (even more so than usual) because we only had access to the water remaining in the loft tank until we could sort this out.

It’s a sobering reminder of the hostility of our surroundings when something goes awry that threatens our Martian livelihood. Admittedly, we have slowly slipped into complacency over the past few days, and this was a stark wake-up call. We troubleshooted as much as we could without going on EVA and came up with a few theories for the malfunction. We decided to give it some time and minimize our aqua consumption in the meantime.

It’s also important for us to remember that EVA is a privilege, rather than a right. Our main goal is to survive on this deceptively deadly planet – the work of thousands back on Earth rides on our perseverance.

In light of this, the crew has taken our lack of EVA-ability in stride! We spent a lot of time today planning our GreenHab reshuffle and did some preliminary reorganization of the plants. We only made the urgent changes today – namely moving the shriveling sprouts out of the path of the heater. That unit throws out heat to the point where our snow peas were starting to look like the plant equivalent of the guy who opens the Ark of the Covenant in Raiders of the Lost Ark. Now they are happily hanging above the spices. Hopefully they turn around and recover from being blasted by the industrial-size hairdryer!

We may have committed an OSHA violation or two while working in the GreenHab, but that’s alright – there’s no rules on Mars! (Sarcasm – we are using our best judgment and staying safe, don’t worry!) However that begs the question: are there laws on Mars? If there are, it probably adheres to international maritime law. That means our Hab is technically a naval vessel! I must spend some time over the next few days coming up with a good nautical name for our trusty ship. But I digress.

Once we headed back to the Hab, we tested the pump once more and it worked with flying colors. Turns out the pipes had frozen overnight – all we needed was a little patience and a lot of sun to remedy our drought. Those little rays traveled millions upon millions of miles from the center of our solar system just to reestablish our access to drinking water. We are eternally grateful. Now to do all those dishes that we let accumulate…

The biggest bright side of our time indoors is it offers us the opportunity for heaps of crew bonding over chores, movies, and games. Synergy is critical for efficient operations in any context, and we are becoming a tight-knit bunch. When we do receive a window for EVA, we are going to be the most productive, well-oiled machine of a crew on the entire planet. In hindsight, I suppose that’s a pretty easy benchmark to reach seeing as we’re the only people on the planet… but nonetheless I’m confident that we’ll exceed expectations.

And not a moment too soon – the meteorological powers that be have predicted quite a few consecutive days of sun. Just the medicine we need to clear up the white blanket covering our landscape. If all goes well, we may be lucky enough to get some EVA time on Sunday! Regardless, priority number one remains the same: survive. We conquer this unforgiving Martian frontier to improve life for those on Earth.

Operations Report – December 27th

Crew 218 Operations Report 27-Dec-19

SOL: 05
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 13
From what time last night: 1930
To what time this morning: 0830
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night: 76 %
Diesel Reading – 83%
Station Propane Reading – 78%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 43
Water Meter (units): 0147061.8
Water (static tank) (gal): 410
Static to Loft Pump used – Yes
Water in Green Hab (gal): 110
Water in Science Dome (gal): 0
Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes
Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? No
CrewCar used and why, where? No
General notes and comments: N/A
Summary of the internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: The pipe from the static tank through the loft tank frozen. The crew rationed the water during the morning until 3:00 PM that the water in the pipes melted. The water system from the Static tank to the Loft tank is working as usual.
Summary of GreenHab operations: routine watering. Moved hanging plants: snow peas and buttercrunch away from heater vent. They are now hanging over the herbs. Recommend transplanting herbs to pots, and transplanting larger plants (tomatoes, peppers, etc.) to Plot E.

Summary of Science Dome operations: Nominal
Summary of RAM operations: Spare fire alarm in the RAM storage closet went off at 12.20 PM and 12.40 PM. HSO removed the battery.
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns, and requests to Mission Support: NA

Luz Ma

ENG

Crew 218

Sol Summary – December 27th

Sol: 05

Summary Title: Still Waitin’ On A Sunny Day

Author’s name: Pat Pesa

Mission Status: Crew is excited for the good weather to come!

Sol Activity Summary: Crew 218 is finding new tasks to keep us involved despite the lack of EVAs due to weather. Today some improvements were made to the GreenHab, and more to still be done tomorrow.

Work also continued with the Engineer’s Indoor Climate-Comfort project.

Look ahead plan: No EVA planned for tomorrow, due to the weather conditions, but sunny and dry days are on the horizon! Future GreenHab improvements will take place pending approval, and the Crew is planning on creating facility training documents/videos for future crews.

Anomalies in work: None

Weather: Intermittent flurries, around 30 degrees

Crew physical status: Healthy

EVA: none

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, HSO incident report

Support Requested: none

Pat Pesa

Geologist, MDRS 218

GreenHab Report – December 27th

Crew 218 GreenHab Report 27-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 22.7°C; 18%

11:40AM

Floor Unit: 16°C

Electronic: 20.7°C

humidity 20%

3:45PM

Floor Unit: 17°C

Electronic: 22.5°C

humidity 19%

5:45PM

Floor Unit: 15°C

Electronic: 20°C

humidity 20%

Max: 27.3°C; 21%

Min: 18°C; 16%

Hours of supplemental light: Light system 7:20PM-12:20AM.

Daily water usage for crops: 6.4Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~110 Gal.

Time(s) of watering for crops:

3:45PM 2.1Gal. watered all hanging plants.

5:45PM 4.3Gal. watered cucumber spinach lakeside, green leaf chard, and wild rocket twice—soil moisture was already dry from earlier watering.

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

—Moved hanging plants: snow peas and butter crunch away from heater vent. They are now hanging over the herbs.

—Recommend transplanting herbs to pots, and transplanting larger plants (tomatoes, peppers, etc.) to Plot E.

Harvest: N/A

Commander Report – December 27th

Crew 218 Commander Report 27-12-2019

Sol 5 – What would a Martian do?
The forecast we received from mission support yesterday proved correct and when we woke up, we could see snow coming down again and covering (this time with a smaller thickness) the terrain that was beginning to show up. Since this means at least one or two more days without EVAs, I called a meeting with the crew after breakfast, to ask ourselves the question "what would a Martian do?". In particular, we all know we are here to do something valuable, besides the obvious chance of studying ourselves in these conditions of isolation Knowing that many of our research projects are based on outdoor activities, we discussed our goals for the next couple of days. In the meanwhile, we waited for the pipes to thaw in order to be able to pump more water inside the habitat, which finally we were able to do in the early afternoon. For our extra goals, we had various proposals that I and other crew members brought up: GreenHab projects, outreach, and use of the solar observatory and science dome were talked about. We began with a simple task that occupied less for about one hour in the GreenHab (more details in other reports), and we discussed a larger project that will be proposed to the program director; we decided that we will produce descriptive and training videos of our life at MDRS; and we will wait for the Sun to finally reappear for some good astronomical observation. I think this was a good opportunity to figure out how to react, with our limited resources, to unexpected situations.

Cesare Guariniello, Commander

Science Report – December 26th

Crew 218 Science Report 26-DEC-2019
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

Nothing to report.

2. Move the first steps towards autonomously generating reliable weather information in the area surrounding MDRS:

Nothing to report.

3. Subsurface structure on Mars:

Nothing to report.

4. Detecting radio signal strength:

Nothing to report.

5. EVA workload analysis:

Nothing to report.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

Nothing to report.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Discussed the progress of measuring environmental stressors inside the habitat with Crew Engineer Luz Maria Agudelo Urrego. These first couple days were calibrating 3 different instruments within her room. She plans to take nighttime measurements tonight. The goal is to develop a continuous dataset, and translate the temperature-humidity combinations into known human thermal comfort algorithms.

8. Messier and other space objects for outreach:

Nothing to report.

9. Reliability and maintenance:

Nothing to report.

10. Medical readings in preparation for future crew-wide project:

Nothing to report.

11. Collection of clay, shale, and hematite samples:

Nothing to report.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.